### v2

# LCDB 1.1: A Database Illustrating Learning Curves Are More Ill-Behaved Than Previously Thought

Abstract
Sample-wise learning curves plot performance versus training set size. They are useful for studying scaling laws and speeding up hyperparameter tuning and model selection. Learning curves are often assumed to be well-behaved: monotone (i.e. improving with more data) and convex. By constructing the Learning Curves Database 1.1 (LCDB 1.1), a large-scale database with high-resolution learning curves including more modern learners (CatBoost, TabNet, RealMLP, and TabPFN), we show that learning curves are less often well-behaved than previously thought. Using statistically rigorous methods, we observe significant ill-behavior in approximately 15% of the learning curves, almost twice as much as in previous estimates. We also identify which learners are to blame and show that specific learners are more ill-behaved than others. Additionally, we demonstrate that different feature scalings rarely resolve ill-behavior. We evaluate the impact of ill-behavior on downstream tasks, such as learning curve fitting and model selection, and find it poses significant challenges, underscoring the relevance and potential of LCDB 1.1 as a challenging benchmark for future research.
1 Introduction
In machine learning, a learning curve can refer to two types of curves. The epoch-wise learning curve (also known as training curve), depicts model performance versus training iterations or epochs. The sample-wise learning curve focuses on performance versus the amount of training data used for training [1]. Sample-wise curves provide a richer evaluation at multiple training data sizes [2]. These curves are useful for speeding up model selection and hyperparameter tuning using multi-fidelity techniques [3]. The curves are also useful to estimate how much data is needed to reach a particular performance [4, 5], providing insights into so-called scaling laws [6, 7, 8, 9]. In this work, we focus exclusively on sample-wise learning curves and use the term learning curve to refer to them.
To effectively use learning curves, it is important to know their shape. When a suitable parametric formula can be assumed, it becomes possible to extrapolate the final performance from partial training data, thereby accelerating model selection. However, learning curve modeling remains challenging: existing parametric models often fail to outperform the simple strategy of selecting the best algorithm based on the last observed curve value [10, 11]. Furthermore, much remains unknown about the learning curve shape. Often, it is assumed that more data leads to better generalization performance [1, 10]. Such learning curves are called monotone: the loss decreases monotonically with more data. Similarly, learning curves are often assumed to be convex, meaning that there is an effect of diminishing returns: the more data we have, the less performance is improved by additional data. If a curve is monotone and convex, we call it well-behaved [12, 13, 1].
Table 1: Ill-behaved learning curves in the wild on OpenML CC-18 classification datasets. The y-axis indicates the error rate on the validation set, and the x-axis represents the size of the training set. The line is the mean and the shaded area indicates one standard error; this applies to all subsequent learning curve plots. Peaking: error rate has a local maximum. Dipping: error rate worsens and does not recover. Phase transition: sudden improvement.
Peaking Dipping Phase Transition
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
non-monotone
non-convex
non-monotone
non-convex
non-monotone
convex
non-monotone
non-convex
monotone
non-convex
Learning curves can exhibit a variety of ill-behaved shapes in toy-settings, violating either monotonicity or convexity [14, 1]. Mohr et al. [10] studied whether such behaviors occur in non-toy settings. They collected the largest-scale database of learning curves, called the Learning Curves Database 1.0 (LCDB 1.0), from various learners, evaluated across hundreds of classification datasets [15, 16]. In LCDB 1.0, Mohr et al. [10] conclude: “We found that the large majority of learning curves is, largely, well-behaved, in that they are monotone, convex, and do not show peaking.” We believe, however, that this conclusion is premature. Their analysis only shows that more extreme ill-behaviors are less frequent; however, it does not estimate how many are significant as a fraction of all curves.
The LCDB 1.0 also suffers from technical issues. It lacks resolution, which can make it difficult to find ill-behavior reliably. In LCDB 1.0, features were also not scaled. Feature scaling is a well-established and standard practice in machine learning that improves training stability and model performance [17, 18, 19, 20]. Therefore one may also wonder if the ill-behavior may disappear simply by scaling. Indeed, we find that sometimes feature scaling makes a curve well-behaved, see Figure 1. Besides, we find that LCDB 1.0 suffers from missing data and a minor data-leakage issue. These issues illustrate the need for a new database and deeper analysis of the prevalence of ill-behaviors.
We introduce the Learning Curves Database 1.1 (LCDB 1.1) which addresses the aforementioned limitations. We incorporate four-times more training set sizes, see Figure 1. This increases the resolution of the curves, which allows us to find more ill-behavior. We argue that, depending on how learning curves are used, data-leakage is sometimes acceptable. Therefore, we introduce two database versions, one with data-leakage and the other without, and also incorporate different feature scalings. In case a performance value is missing due to an error, we justify and document it. Besides, we include the OpenML CC-18 datasets [21], which are more carefully curated datasets, and some more modern tabular data learners, including boosting (CatBoost [22]), deep learning (TabNet [23], RealMLP [24]), and foundation models (TabPFN v2 [25, 26]).
Refer to caption(a)
Refer to caption(b)
Figure 1: Motivation for new LCDB 1.1 features. (a) Feature scaling can mitigate an ill-behaved learning curve. (b) Low-resolution curves may omit certain phenomena or render them less apparent.
Next to providing a new database, we provide a richer analysis of the ill-behaved learning curves. We develop methods to detect whether a learning curve is significantly non-monotone or non-convex and also measure the size of violations. Besides, we also identify other learning phenomena, such as peaking, dipping, and phase transitions, see Table 1. We demonstrate that these ill-behavior are significant and happen often for particular learners. Feature scaling cannot mitigate these ill-behaviors in most of the cases, ruling out that these issues can be easily resolved.
So, we show that these ill-behaviors are significant, but are they also relevant for downstream tasks such as learning curve fitting and model selection? We conduct learning curve fitting experiments using parametric formulas. We investigate the relation between curve fitting and ill-behavior. Most parametric curve models lack the flexibility to model non-monotone and non-convex behavior [27], and indeed we confirm that parametric modeling is significantly more difficult on ill-behaving curves. Learning curves can be used for multi-fidelity model selection using Successive Halving (SH). We find that the crossing curves in our database also make multi-fidelity model selection more challenging. Thus, we illustrate the relevance of these ill-behaviors for downstream tasks, and the unique challenges posed by sample-wise learning curves.
In summary, we create a new and improved database. We perform a more in-depth analysis regarding ill-behaved curves. These analyses illustrate what is inside our database, and therefore we call these database analyses (DA), that we pose as questions. DA1: How many learning curves are significantly ill-behaved and which learners are responsible? DA2: How does feature scaling affect ill-behavior? DA3: How do ill-behaved curves affect learning curve extrapolation? DA4: How does learning curve crossing affect model selection using successive halving?
We discuss the design of LCDB 1.1 in Section 3. Section 4 describes how to robustly detect ill-behaviors. The experimental setup and the results are in Section 5, and we end with discussion and conclusion in Section 6. First, we cover preliminaries and related work.
2 Preliminaries and Related Work
Sample-Wise Learning Curves: Theory and Practice.
A sample-wise learning curve returns a performance $C(n)$ versus the training set sizes $n$ used. Here we discuss its theoretical definition and how it is computed in practice, focusing on classification tasks. Let $S_{n}$ be the training set, consisting of features $x\in\mathbb{R}^{d}$ and corresponding class labels $y$, thus $S_{n}=\{(x_{1},y_{1}),\ldots,(x_{n},y_{n})\}$. We assume that there exists a distribution $P$ from which $(x,y)$ are independent and identically distributed samples. $A(S_{n})$ is a learning algorithm trained on $S_{n}$. Let $R(A(S_{n}))$ be the risk, which indicates its loss in expectation on data from $P$. For classification commonly the zero-one loss is used, in this case, $R(A(S_{n}))$ is the error rate. The theoretical mean learning curve $C(n)$ is defined as:
$C(n)=\mathbb{E}_{S_{n}\sim P}R(A(S_{n})),$ (1)
The curve is computed over a number of training set sizes, e.g. $n_{1},n_{2},\ldots$, where we call the training set size anchor. The risk is an expectation that relies on an integral over the true data distribution $P$, which is unknown. Therefore, we estimate the risk using performance on held out data (test data). The expectation over $S_{n}$ is approximated using multiple repeats with different train and test sets. By using multiple repeats $K$, we obtain multiple estimates of the risk, $\hat{R}_{n}^{r}$, where $n$ indicates the training set size and $r$ is the repetition. We estimate the mean learning curve as $\hat{C}(n)=\frac{1}{K}\sum_{r=1}^{K}\hat{R}_{n}^{r}$. One decides $n_{1},n_{2},\ldots,n_{N}$, typically based on the dataset size. $N$ is the amount of anchors in a curve.
Ill-Behaved, Non-Monotone, and Non-Convex Learning Curves.
Several synthetic learning problems illustrate that more data does not lead to better performance [28, 14], we find such examples in carefully curated CC-18 datasets [21] (see Table  1). Peaking is such a violation, where the error rate initially decreases as the training set size increases, then rises to a peak before decreasing again [29, 30], and is also called sample-wise double descent [31, 5]. Peaking has been proven to occur for the Fisher classifier, and the peak effect is most severe when the training set size is equal to the dimensionality $d$ [32, 33]. Double descent describes a similar phenomenon when plotting the error versus the capacity of the model [34, 1]. Other cases of monotonicity and convexity violations include dipping and phase transitions. Loog and Duin [35] introduce the concept of dipping. The error rate initially improves, then increases without recovering, even in the limit of infinite amounts of data. Dipping has been observed for decision trees in error rate learning curves [36, 37], and for KNN in AUC learning curves [38]. A phase transition means that model performance improves abruptly, causing a distinct drop in the learning curve. Phase transitions in machine learning were studied mostly in theory [39, 1], and we are not aware of any examples on real-world datasets before this work. Beyond classification, many of the observed learning curve irregularities, such as non-monotonicity, may also arise in regression problems [40, 41, 14] and even unsupervised learning [42], challenging the assumption that more data always helps.
The Learning Curves Database 1.0 (LCDB 1.0).
The LCDB 1.0 [10] includes classification learning curves of various learners on numerous datasets from the OpenML platform [43, 44, 45]. Some curves in LCDB 1.0 are missing, resulting in actually fewer than 196 datasets, possibly due to incompatibilities with sparse matrices and long compute times. Data leakage occurs because the feature imputer was fitted on the complete data. We resolve these issues with the LCDB 1.1.
Other Datasets and Relation to Deep Learning.
Task-Set [46], LCBench [47], and BUTTER [48] are datasets containing epoch-wise curves of neural networks. They are not comparable to ours, since we study sample-wise learning curves. In deep learning, the scaling law literature focuses on much larger training sets and presents much sparser learning curves [7, 8, 9]. LCDB 1.1 instead focuses on tabular data, where many classical algorithms remain competitive with deep learning [49, 50, 51, 52], and where datasets are typically smaller. Furthermore, tabular data offer unique challenges: these datasets are rare [53], and columns are often incomparable across datasets, complicating knowledge transfer [54, 55]. Meanwhile, tabular data are crucially important for industry [56]. Insights into learning curves can help estimate how much tabular data is needed [5] which is important when data is costly.
Learning Curve Fitting.
Learning curves are usually modeled by parametric formulas [1, 8, 57]. Popular functions are exponential and power laws, which are motivated by the well-behaved assumption [1]. Learning curve fitting can be used to estimate the amount of data needed [5]. Mohr et al. [10] identified that parametric models with 4 parameters seem to perform best for interpolation, such as $\mathrm{POW4}$, where $\hat{C}(n)=a-b(d+n)^{-c}$. The most widespread technique is least square curve-fitting using Levenberg-Marquadt [28], more advanced techniques use Bayesian techniques and neural networks [58, 59, 60]. We investigate the effect of ill-behavior on least square curve fitting.
Multi-Fidelity Model Selection.
Successive Halving [61] (SH) is a method to speed up model selection. It uses a fidelity; the higher the fidelity, the more accurate model performance is estimated. The fidelity can represent the amount of epochs used or the amount of training data. SH is iterative, evaluating model performances first at low fidelities and moving to higher fidelities afterward. In each round, a percentage of the learners that perform poorest are dropped, and the fidelity is increased. SH can be combined with learning curve extrapolation for both learning curves [13] and training curves [58]. If learning curves often cross, SH may perform suboptimal, which we will investigate. Various multi-fidelity methods exist [62, 58, 63, 64, 65, 66, 67], we use SH since it is popular and interpretable.
3 The Improved Learning Curves Database 1.1
Table 2 gives an overview of the main differences between LCDB 1.1 and 1.0. First, we discuss data splitting, preprocessing and we justify two the versions with and without data-leakage. We briefly discuss dataset and learner selection, and end with metrics, reproducibility, and some statistics. Some details equal to LCDB 1.0 are omitted (see Appendix A). The LCDB 1.1 is publicly available. [^1] LCDB 1.1 dataset: https://doi.org/10.4121/3bd18108-fad0-4e4c-affd-4341fba99306
Table 2: Main innovations of LCDB 1.1 compared to LCDB 1.0.
Database Preprocessor Feature Scaling Anchor Resolution #Learners #Datasets Missing Missing Reason
LCDB 1.0 with data-leakage (dl) none $\lceil 16\cdot 2^{k/2}\rceil$ 20 196 (claim 246) 12% (30%) unknown
LCDB 1.1 with and without dl none, min-max, standard 4 times denser 32 265 4% documented
Data Splitting.
We use multiple train-validation-test sets to enable the simulation of model selection using nested cross validation. Selection can be done using validation and evaluation using the test set. We use 5 inner and 5 outer seeds to create these datasets. Let $D$ be the complete dataset, then
$D\xrightarrow[\text{outer seed}~m]{\text{outer split}}\left(D_{\text{train-val}}^{(m_{o})},D_{\text{test}}^{(m_{o})}\right)\quad\text{then}\quad D_{\text{train-val}}^{(m)}\xrightarrow[n\,\text{random seed}]{\text{inner split}}\left(D_{\text{train}}^{(m_{o},m_{i})},D_{\text{val}}^{(m_{o},m_{i})}\right)$
where the superscripts indicate outer ($m_{o}$) and inner ($m_{i}$) seeds. LCDB 1.0 uses training anchors $n_{k}=\lceil 16\cdot 2^{k/2}\rceil$, where $k\in\{0,1,2,$...\}. The LCDB 1.1 uses $n_{k}=\lceil 16\cdot 2^{k/8}\rceil$ resulting in four times higher resolution. Further details are as in LCDB 1.0 (see Appendix A).
On Preprocessing and Data Leakage.
In LCDB 1.0, the imputer was fitted on the whole dataset. In LCDB 1.1, we apply: no scaling (abbreviated as “noFS”), min-max scaling, or standardization of features. Because of this additional preprocessing, it is even more important to discuss how to fit the preprocessor and data-leakage. When learning curves are applied for model selection and hyperparameter tuning, the goal is to reduce computation time. We can assume the user has access to the complete dataset. Fitting the preprocessor on the whole dataset can then lead to better performance and stability, and data-leakage is acceptable. However, when trying to estimate how much data is needed, we cannot assume the user has access to all data. Thus, in this case, data-leakage is not acceptable. We therefore construct two LCDB 1.1 variants, with and without data-leakage. To prevent data-leakage, preprocessors are fitted on the train set. We compare these versions in Appendix B.
Dataset and Learner Selection.
In LCDB 1.0, we observe that some datasets are overly easy, resulting in flat learning curves that are already converged at the first anchor. Therefore, we include all datasets of the OpenML-CC18 benchmark [21] in LCDB 1.1, called LCDB 1.1 CC-18. This benchmark was carefully curated, filtering out datasets that are overly easy, amongst other issues. The complete LCDB 1.1, referred to as LCDB 1.1 FULL, combines CC-18 with all datasets of LCDB 1.0.
The LCDB 1.1 has 32 learners, see Table 3. The dummy predicts the majority class and provides a weak baseline. One-hot features violate assumptions of Naive Bayes [68], to that end we introduce mixed Naive Bayes learners. Moreover, we incorporate a broader set of modern tabular learners: the boosting model CatBoost [22], deep learning models such as TabNet [23] and RealMLP [24], and the foundation model TabPFN v2 [26]. According to Erickson et al. [69], CatBoost remains a strong state-of-the-art model by default, while RealMLP achieves state-of-the-art performance after tuning and ensembling. TabNet is a popular deep learning baseline [23], and TabPFN v2 is a well-performing foundation model for tabular data [26]. See Appendix A for all added learners.
All these modern learners claim robustness to differently scaled features; therefore, we only include their no scaling (noFS) and no data-leakage variants in LCDB 1.1 FULL (except TabPFN, which does not explicitly address feature scaling) [22, 23, 24, 26]. Regarding categorical features, RealMLP and TabPFN use one-hot encoding, whereas for CatBoost and TabNet we follow their suggested practice of directly feeding categorical features into the model. For implementation details see Appendix A.
Table 3: The 32 learners in LCDB 1.1 FULL (265 OpenML datasets, no scaling version), their estimated ill-behaved (non-monotone or non-convex) ratio, and their abbreviations.
Learners (Abbreviation) Ill-behaved Learners (Abbreviation) Ill-behaved
CatBoost [22] 1.5% Complement Naive Bayes (ComplementNB) 8.3%
Decision Tree (DT) 1.5% Passive Aggressive (PA) 9.4%
TabPFN v2 [26] * 1.5% Mix Complement Naive Bayes (MixComplementNB) 10.2%
Extra Tree (ET) 1.9% Mix Multinomial Naive Bayes (MixMultinomialNB) 10.6%
ensemble Gradient Boosting (ens. GB) 1.9% RBF Support Vector Machine (SVM_RBF) 15.8%
ensemble Random Forest (ens. RF) 3.0% Ridge Regression Classifier (Ridge) 17.0%
Stochastic Gradient Descent Classifier (SGD) 3.4% Mix Gaussian Naive Bayes (MixGaussianNB) 21.5%
ensemble Extra Trees (ens. ET) 3.4% Gaussian Naive Bayes (GaussianNB) 24.9%
Perceptron 3.8% Multilayer Perceptron (MLP) 27.9%
K-Nearest Neighbors (KNN) 3.8% Bernoulli Naive Bayes (BernoulliNB) 28.3%
RealMLP [24] ** 5.3% Mix Bernoulli Naive Bayes (Mix BernoulliNB) 28.7%
Logistic Regression (LR) 5.3% Linear Discriminant Analysis (LDA) 37.7%
Linear Support Vector Machine (SVM_Linear) 5.7% Quadratic Discriminant Analysis (QDA) 45.7%
Polynomial Support Vector Machine (SVM_Poly) 7.9% Sigmoid Support Vector Machine (SVM_Sigmoid) 58.1%
Multinomial Naive Bayes (MultinomialNB) 7.9% Dummy Classifier (Dummy) 60%
Nearest Centroid (NC) 7.9% TabNet [23] 74.3%
    The reported results cover 210 out of 265 datasets with the maximum curves length less than 10k, due to the fact that TabPFN v2 only supports datasets with up to 10k training samples, 500 features, and 10 classes. Note that some datasets included in LCDB 1.1 were used in designing its prior.
    Some LCDB 1.1 datasets were used in the meta-train benchmark for designing and meta-tuning RealMLP.
Metrics, Reproducibility, and Database statistics.
We use Python, scikit-learn [70], a docker image, save all package versions and provide all the code for reproducibility. [^2] https://github.com/learning-curve-research/LCDB-1.1 We fix the seed of the learner to make them reproducible. We compute: error rate, F1, AUC, and log-loss for the validation and test sets, and we store learners’ scores or probabilistic outputs. When training fails, we record the error message and set the performance to Not-A-Number (NaN). Table 4 shows the proportions of different curve shapes (their detections are discussed in the next section). Missingness refers to NaN values and is mostly caused by Naive Bayes learners that cannot handle negative features. While not the main point of this work, one may wonder which learners perform best; see Appendix C.
4 Robustly Measuring Monotonicity, Convexity, Peaking and Dipping
In this section, we introduce the methods to detect and also measure monotonicity violations, convexity violations, peaking and dipping curves. However, we first criticize the methods of Mohr et al. [10]. They analyze monotonicity and convexity, but only consecutive anchors are compared. This will miss violations that happen over longer ranges, which is why we compare all anchors. We also check for significance, and since neighboring training set sizes may not yield significant differences, comparing all pairs is even more crucial. The convexity measure of Mohr et al. [10] treats the anchors as linearly spaced, ignoring that they are defined in logarithmic scale, which leads to incorrect conclusions. Our method incorporates the anchor scale. We use a hypothesis test to ensure detections are significant, where we are pessimistic, e.g. we only find violations if we are confident, otherwise we assume the curve is well-behaved. This aligns with the prior belief that most curves are well-behaved following literature [1, 10]. We assume a metric $C$ where lower means better.
A monotonicity violation means that the curve does not always improve with more data, see Figure 2.
Definition 1 (Monotonicity Violation Error).
The largest increase between any anchor pair is
$\epsilon_{\text{mono}}=\max\left(0,\max_{1\leq i<j\leq N}\left(C(n_{j})-C(n_{i})\right)\right).$ (2)
The violation error $\epsilon_{\text{mono}}$ measures the largest size of the violation and is zero if there is none.
Figure 2: Monotonicity Violation
To detect violations from empirical learning curves, we use the following procedure. We compute $\hat{\epsilon}_{mono}$ using the empirical curve means and find the pair $(i^{*},j^{*})$ that maximizes Equation 2. If $\hat{\epsilon}_{mono}$ is zero, we classify the curve as monotone. If $\hat{\epsilon}_{mono}>0$ we check the significance of the violation. We compare the empirical distributions $\hat{R}^{r}_{n_{i^{*}}}$ and $\hat{R}^{r}_{n_{j^{*}}}$ using a paired one-sided t-test with Bonferroni correction. Paired, because the same inner and outer seeds are used, and one-sided because we only care about violations in one direction. The Bonferroni correction corrects for multiple testing, assuring we do not find too many violations due to noise. This correction is necessary because identifying the maximum among anchor pairs implicitly involves multiple comparisons. We correct on a curve-level for all anchor pairs. The corrected significance level is $\alpha^{\prime}=\frac{\alpha}{N(N-1)/2}$, where $\alpha$ is the original significance level. If the $p$-value is smaller than $\alpha^{\prime}$ we classify the curve as non-monotone.
A function is convex if its linear interpolation is always above the function itself. If the curve is above its linear interpolation, this is a convexity violation, see Figure 3.
Definition 2 (Convexity Violation Error).
The linear interpolation of a curve from anchor $n_{h}$ to $n_{j}$ evaluated at $n_{i}$ is: $C_{\text{interpolated}}(n_{i};n_{h},n_{j})=\frac{n_{j}-n_{i}}{n_{j}-n_{h}}\,C(n_{h})+\frac{n_{i}-n_{h}}{n_{j}-n_{h}}\,C(n_{j})$. We define
$\epsilon_{conv}=\max\left(0,\max_{1\leq h<i<j\leq N}\left(C(n_{i})-C_{interpolated}(n_{i};n_{h},n_{j})\right)\right).$ (3)
The violation error $\epsilon_{conv}$ measures the largest convexity violation and is zero if there is none.
Figure 3: Convexity Violation
We detect a convexity violation from empirical data using the following procedure. First, we evaluate $\hat{\epsilon}_{conv}$ using the empirical means of the learning curve. If $\hat{\epsilon}_{conv}<0$, the curve is classified as convex. If $\hat{\epsilon}_{conv}>0$, we check the significance of the violation. First, we find the maximizers $(i^{*},j^{*},h^{*})$ of Equation 3. For each repeat, we interpolate the curve linearly, to obtain the empirical distribution of the interpolated curve. The interpolated and actual distributions are compared using one-sided paired $t$-test. We correct for the triplet comparison using Bonferroni; thus if the $p$-value is smaller than $\alpha^{\prime}=\frac{\alpha}{N(N-1)(N-2)/(3!)}$, we classify the curve as non-convex.
Definition 3 (Peaking Phenomenon).
Peaking occurs if there exists a triplet of indices, $1\leq h<i<j\leq N$, such that:
$C(n_{i})>C(n_{h})\quad\text{and}\quad C(n_{i})>C(n_{j}).$ (4)
In this case, $C(n_{i})$ forms a local peak, indicating that the model’s performance temporarily degrades and subsequently recovers as more data is added.
Definition 4 (Dipping Phenomenon).
Dipping occurs if there exists an index $i$, $1\leq i<N$, such that:
$C(n_{i})<C(n_{N}).$ (5)
NN denotes the amount of anchors in a curve. This indicates a sustained degradation of model performance, with no recovery observed as more data is added.
Lastly, we describe how peaking and dipping are detected. Peaking is characterized by a combination of convexity and monotonicity violations: we first locate a convexity violation at $(h^{*},i^{*},j^{*})$, and then verify a monotonicity violation between $h^{*}$ and $i^{*}$ and we check for significant improvement between $i^{*}$ and $j^{*}$ (similar to violation error detection, but instead checking for improvement). If all 3 conditions are satisfied, the curve is classified as peaking. Dipping corresponds to a monotonicity violation with $j$ fixed as the last anchor $N$.
5 Results
Here we discuss the database analyses (DA) that we perform and the experimental setup.
5.1 Experimental Setup
Both QDA and the Dummy classifier are excluded due to reproducibility issues and the lack of meaningful learning behavior, respectively. We do not conduct analyses of mixed Naive Bayes methods, as their curves are largely indistinguishable from standard Naive Bayes (Appendix C). We always focus on error rate learning curves in LCDB 1.1 CC-18, since its selection of datasets is more carefully curated. Results on LCDB 1.1 FULL are similar (see Appendix D). DA1 and DA2. A significance level of $\alpha=0.05$ is used throughout, and the curves are estimated using the validation set. Since we have 5 inner and 5 outer seeds, the learning curves are estimated from 25 repeats, which are aggregated together. To compare with LCDB 1.0, we interpolate the LCDB 1.0 curves to have the same length to ensure Bonferroni corrections are comparable. DA3. We closely follow the curve fitting methodology of LCDB 1.0 [10] and also use the validation set. We use the parametric models $\mathrm{POW4}$, $\mathrm{MMF4}$, and $\mathrm{WBL4}$ since they performed best. Flat curves are filtered because they are overly easy to fit, leading to very small MSEs. To detect them, we scale all learners’ curves to [0,1] range and classify it as flat if the maximum minus minimum value is below 0.05. DA4. We run successive halving to perform model selection, where the fidelity is determined by the anchor. Model selection is done using the validation set, and the selected model is evaluated using the test set.
5.2 DA1: How Many Curves Are Significantly Ill-Behaved and Which Learners Are Responsible?
An overall picture of the violations can be observed in Table 3 and 4. In this section, we only discuss the no feature scaling case (“no FS”). A substantial amount of curves is non-monotone (9.9%) and non-convex (11.5%), leading to 14.9% ill-behaved curves. This is significantly larger than the significance level $\alpha$, ruling out that these curves ill-behaviors are purely caused by noise. Note that the LCDB 1.0 barely passes this bar, underlining the need for a higher resolution database like LCDB 1.1 to detect all ill-behaviors. Peaking is responsible for 5.0% and dipping is responsible for 6.1%. The amount of flat curves is reduced for the CC-18 version compared to the FULL version as expected due to more careful curation.
Table 4: Ill-behavior statistics of the LCDB 1.1 variants and LCDB 1.0. Since we use a significance level of $5\%$ to detect ill-behaviors, we can expect $5\%$ false positives (in the worst-case). Therefore, only numbers larger than $5\%$ are significant, which the LCDB 1.0 barely satisfies. Note: “no FS” results include the statistics of 4 more modern learners.
Shapes / Database LCDB 1.1 CC-18 (72) LCDB 1.1 FULL (265) LCDB 1.0 (196)
no FS min-max FS standardization FS no FS min-max FS standardization FS no FS with interp.
Missing 2.1% 0.0% 8.7% 3.0% 0.0% 8.7% 11.9%
Flat 7.1% 5.8% 3.4% 9.9% 7.9% 5.3% 5.2%
Non-Monotone ($\neg$ M) 9.9% 11.2% 9.2% 9.6% 11.1% 9.5% 5.1%
Non-Convex ($\neg$ C) 11.5% 9.4% 8.4% 12.3% 10.0% 8.8% 5.7%
Ill-behaved ($\neg$ M $\cup$ $\neg$ C) 14.9% 13.5% 11.2% 15.4% 14.3% 11.8% 8.1%
Peaking 5.0% 3.3% 2.9% 5.7% 3.7% 3.2% 2.5%
Dipping 6.1% 8.5% 6.3% 6.9% 9.6% 7.2% 4.6%
In Figure 4, we visualize the ill-behaviors per learner. Learners that have less than $5\%$ of any of the ill-behaviors are omitted, for a full overview see Appendix D. Again, we discuss the case of no feature scaling. The MLP can exhibit surprising learning curve shapes that we classify as phase transitions (see examples in Appendix E.1). Additionally, we observe several peaking caused by artifacts arising from the interplay between batch size and training set size (see also Appendix E.1). The Sigmoid SVM is a notably ill-behaved learner, showing many monotonicity and convexity violations, of which most can be classified as dipping. The RBF SVM is more well-behaved but does show some peaking. Note that the statistical stringency differs across ill-behaviors. For example, Sigmoid SVM shows more dipping than monotonicity violations, we return to this issue in Section 6.
We also observe peaking for LDA and the Ridge classifier. This can be expected because Ridge and LDA are closely related to Fisher [71, 4.3] [72, 4.1.5] which is known to peak. Surprisingly, KNN, Naive Bayes, and Nearest Centroid also do not always behave well. For Nearest Centroid, it was known it could dip [35], but this was never observed outside of toy settings. It can be concluded that a few learners are in fact responsible for the most ill-behaving curves. The most well-behaved learners are tree-like and ensemble learners (see also Table 3).
While many modern learners, such as CatBoost, RealMLP, and TabPFN v2, tend to exhibit well-behaved learning curves, this is not always the case. The notably ill-behaved learner TabNet exhibits substantial non-convexity in its learning curves (see Appendix D), which mostly appears to stem from phase transition phenomena (see Appendix E.2). We suspect that TabNet was designed for large datasets, and that its default hyperparameters are not suitable for small dataset sizes. Indeed, we find that generally for large training set sizes, TabNet performs well (see Appendix C).
Refer to caption(a)
Refer to caption(b)
Refer to caption(c)
Refer to caption(d)
Figure 4: Estimated probability ($\%$) of different ill-behaviors, (a) Monotonicity Violation, (b) Convexity Violation, (c) Dipping, (d) Peaking, for learners with different feature scalings. For all results see Appendix D. Observe that feature scaling for most learners does not lead to significant changes. Ridge and MLP improve significantly, while NC becomes more ill-behaved.
5.3 DA2: Can Feature Scaling Mitigate Ill-Behavior?
To understand the impact of feature scaling, we now compare the results across scaling techniques. Table 4 indicates that feature scaling marginally reduces the amount of ill-behavior. From Figure 4 we observe that for most learners, feature scaling does not resolve ill-behavior. The Sigmoid SVM becomes slightly more monotone and has fewer peaks, but is still significantly ill-behaved. While preprocessing does not reduce ill-behavior, the SVM absolute performance improves notably and training becomes more stable after scaling is applied (see Appendix E.3 for details). Nearest Centroid and some Naive Bayes models are the only models that become significantly more ill-behaved with feature scaling. Note that GaussianNB is not entirely invariant to feature scaling, due to the way it calculates the variance for numerical stability. The biggest reductions in ill-behavior occur for the Ridge classifier and MLP. Ridge becomes almost completely monotone and without peaks when using min-max scaling, but not with standard scaling. The MLP improves significantly when using standard scaling, largely resolving ill-behaviors, however, min-max scaling does not always help the MLP. We confirm LDA is insensitive to feature scaling, and find the peak occurs when the training set size is approximately equal to the dimensionality (Appendix E.4) in line with peaking literature.
A further analysis showing which datasets are responsible for ill-behavior can be found in Appendix E.5. The ill-behavior seems to occur on almost all datasets, and in particular, it is not possible to attribute ill-behavior to a small number of datasets. In conclusion, few models become more well-behaved with preprocessing, and the type of preprocessing that helps can differ per model.
5.4 DA3: How Do Monotonicity and Convexity Violations Affect Curve Fitting?
In Figure 5 we show how the curve fitting performance is affected by convexity and monotonicity violations. We focus here on the results for the parametric formula $\mathrm{POW4}$ (power law). For MMF4 and WBL4, results are similar, see Appendix F. Performance is measured using the mean squared error (MSE) on the fitted points (interpolation). The mean of the log MSE for monotone curves is over ten times smaller than for non-monotone curves, and the same applies to convex versus non-convex curves. Figure 5 visualizes the MSE versus the violation error. The results reveal a clear positive correlation between violation error and MSE. Our findings clearly show that parametric model fitting is significantly harder for non-monotone and non-convex curves, establishing LCDB 1.1 as a challenging benchmark for evaluating learning curve modeling methods.
Refer to caption(a)
Refer to caption(b)
Figure 5: Ill-behaved learning curves pose challenges for curve fitting. (a) The distribution of fitting MSE when applying a parametric model to monotone vs. non-monotone (left) and convex vs. non-convex (right) learning curves. The dashed lines represent mean of the log MSE. Ill-behavior leads to to significantly larger MSE. (b) Larger violation sizes (x-axis) coincide with larger MSE (y-axis).
5.5 DA4: How Do Crossing Learning Curves Affect Model Selection?
Refer to caption(a)
Refer to caption(b)
Figure 6: (a) Learning curve crossing probabilities of the two learner subsets (blue and orange), heatmap indicates probability. (b) Results of Successive Halving (SH) applied to blue and orange learner subsets. Fewer crossings lead to better SH performance. More results are in Appendix G.
Here, we choose two sets of learners and run Successive Halving (SH) on them to perform model selection with the training set size as fidelity, to investigate the influence of crossing curves. We determine $5$-subsets of learners, one set of learners that often cross, and one set of learners who rarely cross; see Figure 6. On both sets of learners we run SH, the results are shown in Figure 6. In the left figure, we show how often the best algorithm is found. However, since the final performance differences of learners may be very similar, we complement this figure with the regrets on the right. Regret is the final error rate of the chosen learner minus the minimum of the final error rate over the learners in the subset (note the log-scale). Results for more settings are given in Appendix G. In the group of learners whose curves rarely cross (blue), the algorithm almost always picks the best or at least runner-up. For learners that frequently cross this is less often the case. The regrets show a similar pattern. As such, we can observe that crossing curves make model selection using SH significantly more challenging, highlighting the usefulness of LCDB 1.1 as a challenging benchmark for evaluating multi-fidelity model selection strategies.
6 Discussion
In contrast to Mohr et al. [10], we do find significant amounts of ill-behavior using our improved LCDB 1.1. While peaking and dipping were previously known for LDA, Ridge, Nearest Centroid, and MLP [1], their occurrence in realistic settings was not established. For the Sigmoid SVM, Naive Bayes, and KNN, it was not known ill-behavior was possible (either in toy or realistic settings). Ensemble methods have very well-behaved curves, yet we observe severe dipping on OpenML dataset 41027 [73]. The causes of these ill-behaviors remain unclear and present a challenging open problem.
The shape analysis is challenging. Note that ill-behavior may change if curves are longer (especially dipping). It is also difficult to maintain statistical rigor and consistency; this is because the Bonferroni correction imposes different levels of stringency, for instance, when testing monotonicity (two anchors) and convexity (three anchors). Bonferroni is also quite pessimistic, and some subjective choices had to be made. For example, peaking can also be detected differently (see Appendix H), yet results are similar. Moreover, we have performed additional analysis using slightly less conservative method called Holm’s Step-Down Procedure (Holm’s method) [18], in the sense that it will reject more null hypotheses, typically resulting in fewer Type II errors. This slightly increases the proportion of ill-behaved cases to 19%, but preserves overall consistency (see Appendix I). An analysis using E-values [74] or controlling the false discovery rate [75] may alleviate inconsistency issues, but such an analysis is non-trivial, going beyond our main point.
We have tried our best to make LCDB 1.1 fully reproducible, by using a docker container and fixed python package versions, yet we find that one LDA curve and several QDA curves are non-reproducible, likely due to numerical non-determinism of the singular value decomposition. For this reason, we exclude the QDA learner from the analysis but include its curves in the database.
The next step is to investigate whether ill-behavior persists under hyperparameter tuning, which we leave for future work, since collecting learning curve data with tuning is computationally expensive (LCDB 1.1 already required 800K CPU hours; see Appendix J). This investigation is particularly relevant for models that may exhibit strong sensitivity to hyperparameter settings, such as TabNet. Moreover, hyperparameters in Scikit have reasonable defaults determined by the community, and as such the ill-behavior observed remain surprising and relevant, especially when persistent to different scalings. Although our empirical analysis is scoped to error rate learning curves for classification, the observed ill-behaviors are not confined to this setting. It is therefore valuable to examine alternative evaluation metrics, such as AUC, F1 score, and log-loss, all of which we provide in LCDB 1.1.
7 Conclusion
In conclusion, we introduce the Learning Curves Database 1.1 (LCDB 1.1). This database is more reproducible, of higher resolution, and has multiple types of preprocessing (with and without data-leakage), as well as more modern learners such as CatBoost, TabNet, RealMLP, and TabPFN, making it a valuable database for the community to study learning curves. Moreover, we carefully study ill-behavior and find that a significant amount, 15%, of the curves exhibit ill-behavior while some learners misbehave more frequently than others. Feature scaling rarely solves this problem and in some cases can make it worse. Lastly, we demonstrate the impact of ill-behavior on downstream tasks, underscoring the practical implications. We hope that LCDB 1.1 facilitates new investigations of ill-behavior and serves as a challenging benchmark to evaluate downstream tasks.
Acknowledgments and Disclosure of Funding
We gratefully acknowledge Jesse Krijthe, Taylan Turan, and the anonymous reviewers for their valuable reviews and insightful suggestions. We also thank Jan van Gemert, Marcel Reinders, David Tax, and Gijs van Tulder for their constructive discussions. We further acknowledge the support of DAIC (Delft AI Cluster) and DelftBlue for providing computational resources. Moreover, we thank the community, particularly OpenML for enabling the sharing of data and Scikit-learn for providing essential machine learning tools.
Appendix A LCDB 1.1 Additional Details
Data Splitting.
We split the data twice, first, the outer split (outer seed) splits off 10% test data. The inner split (inner seed) splits the remainder in a train (90%) and validation (10%) set. Validation and testing set are capped at 5000 samples. We use 5 inner and 5 outer seeds and these splits are stratified. Training sets are further reduced in size to simulate the collection of a learning curve. The training sets are constructed in a monotonic way without stratification, i.e. $S_{1}\subset S_{2}\subset...\subset S_{n}$. This procedure corresponds exactly with how the LCDB 1.0 also was collected [10].
Imputation.
We impute the median for numerical features and the most frequent value for categorical features to deal with missing data. For categorical features, we apply one-hot encoding. This procedure corresponds exactly with how the LCDB 1.0 also was collected [10]. Note that, for LCDB 1.0, if the number of features is very large, features were binarized, which we believe was not a intended preprocessing step. We do not include any binarization.
Justification of Other Additional Learners.
The Complement Naive Bayes learner was introduced to resolve poor assumptions of the Multinomial Naive Bayes classifier, hence we include it [76]. Complement and Multinomial Naive Bayes are intended for text classification, yet few datasets are text datasets. Therefore, we decided to also include Gaussian Naive Bayes, which assumes features are Gaussian, which can be more reasonable for our diversity of datasets. We, however, choose to include all Naive Bayes learners, as they were also included in the LCDB 1.0. The Nearest Centroid classifier is computationally efficient but is known to display ill-behavior in toy settings [35].
Naive Bayes Preprocessing and Mix-Naive Bayes.
Each Naive Bayes model is included twice: as an original and mixed version. In LCDB 1.0, Naive Bayes was trained on all features, including the one-hot encoded features, which we call original. One-hot encoded features violate the core assumption of conditional independence that underlies the Naive Bayes model [68]. The mixed Naive Bayes models categorical and numerical features separately. Categorical Naive Bayes is used for categorical features, and the other model is used on the numerical features (Bernoulli, Multinomial, Complement, Guassian), ensuring that the categorical features are modeled appropriately.
Modern Learner Implementation Details.
We use the official implementations of CatBoost, TabNet, RealMLP, and TabPFN v2 with their default hyperparameters. For TabNet, we employ the small-scale model (TabNet-S) and use the default hyperparameters without early stopping, different from how TabNet was configured in [23], to ensure consistency across all learners.
Appendix B Difference Between LCDB 1.1 versions
To assess whether data leakage meaningfully alters the learning curves, we computed the proportion of instances where a statistically significant difference (based on Bonferroni-corrected t-tests) was observed between results obtained with and without potential leakage. This comparison was performed across three preprocessing configurations: no feature scaling, min-max normalization, and standardization. If there is one anchor significantly different between two curves, we classify it as a different curve. The results are summarized in Table 5. Observe that, for the case of no feature scaling, the amount of different curves is small, because here only imputation was performed. When feature scaling is used, data leakage becomes more pronounced.
Table 5: Percentage of curves with at least one anchor that is significantly different between data leakage and no data leakage version of the LCDB 1.1.
no FS min-max FS standardization FS
LCDB 1.1 CC-18 (72) 1.2% 12.3% 8.3%
LCDB 1.1 FULL (265) 1.9% 15.6% 12.5%
Appendix C Absolute Performance Comparison
In addition to analyzing the shape of the learning curve, we can also compare the performance of different learning algorithms on different feature scaling techniques by using optimal performance points on the learning curve. As a simple supplementary analysis provided in the appendix, Figure 7 presents the best average error rates of different learners on both LCDB 1.1 CC-18 and FULL. Specifically, we extract the minimum error rate from each learning curve and compute the average across datasets for each learner. It is evident that feature scaling has minimal impact on tree-based algorithms, but can significantly improve the performance of many distance-based and iteratively fitted learners.
Refer to caption(a)
Refer to caption(b)
Figure 7: Comparison of learner best performance on average under different feature scaling strategies. (a) LCDB 1.1 CC-18 (72 datasets). (b) LCDB 1.1 FULL (265 datasets).
Furthermore, we can investigate the learners’ optimal performance by filtering datasets according to their characteristics. This analysis can serve as a simple use case that provides further evidence on whether deep learning models outperform tree-based methods under different scale of dataset. Figure 8 compares the learners’ performance between two groups: one where at least one anchor includes more than 10k training samples, and another where all anchors have fewer than 10k training samples. As shown, CatBoost and RealMLP demonstrate consistently strong performance across both groups, while TabNet exhibits competitive performance only on datasets with larger training set sizes.
Refer to caption(a)
Refer to caption(b)
Figure 8: Comparison of learner best performance on average under different feature scaling strategies. (a) LCDB 1.1 FULL (200 datasets with number of samples less than 10k). (b) LCDB 1.1 FULL (65 datasets with number of samples more than 10k).
Motivated by the analysis in Williams [68], which highlights the incorrect assumption of conditional independence in Naive Bayes when applied to one-hot encoded features, we explore mixed Naive Bayes model variants. The one-hot encoded features are not independent and may lead to inaccurate probability estimates and model misspeficiation. To address this, we introduce mixed Naive Bayes models and evaluate their performance using the win-loss-tie framework on the LCDB 1.1 FULL. We compare the performances over all anchors and datasets, and record a win if one method is better than the other and a tie if they achieve the same performance. As shown in Figure 9, mixed Naive Bayes models do not always outperform their vanilla counterparts. The tie cases are primarily due to datasets without categorical features, where no encoding is applied and both models behave identically. In conclusion, the model-misspecification for Naive Bayes does not seem so problematic for the error rate.
Figure 9: The performance comparison between Naive Bayes and mixed Naive Bayes methods by using LCDB 1.1 FULL (265 OpenML datasets).
Appendix D Statistics in LCDB 1.1 per Learner
Similar to Table 4, we show the statistics of different learners in LCDB 1.1 for both CC-18 and FULL with no data-leakage version in Tables 6, 7, and 8. Here, non-monotone, non-convex, and ill-behaved refer to shapes that violate monotonicity, convexity, and either, respectively. In Table 6, we only include the no feature scaling version, since the three considered models can internally handle feature scaling. Note, some datasets from CC-18 or FULL were used in the meta-train benchmark for designing and meta-tuning RealMLP. For Table 7, the relatively high proportion of missing curves is mainly due to the limitations of TabPFN v2, which only supports datasets with up to 10k training samples, 500 features, and 10 classes, as well as curves whose maximum length is below 10k. Moreover, some datasets included in CC-18 and FULL were used during the design of the TabPFN prior. Therefore, the comparison involving TabPFN v2 is not entirely fair, and we present this table separately.
Table 6: Statistics of CatBoost, TabNet, and RealMLP in LCDB 1.1 (no feature scaling).
Learner / Ratio(%) LCDB 1.1 CC-18 (72) LCDB 1.1 FULL (265)
Missing Non-Monotone Non-Convex Ill-behaved Peaking Dipping Missing Non-Monotone Non-Convex Ill-behaved Peaking Dipping
CatBoost 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 0.0% 0.4% 1.1% 1.5% 0.8% 0.4%
TabNet 0.0% 11.1% 72.2% 72.2% 33.3% 1.4% 0.4% 17.4% 73.6% 74.3% 42.3% 4.2%
RealMLP 0.0% 0.0% 1.4% 1.4% 0.0% 0.0% 0.0% 0.4% 4.9% 5.3% 0.0% 0.0%
Table 7: Statistics of TabPFN v2 in LCDB 1.1.
Learner / Ratio(%) LCDB 1.1 CC-18 (72) LCDB 1.1 FULL (265)
no FS min-max FS standardization FS no FS min-max FS standardization FS
Missing 12.5% 12.5% 12.5% 20.8% 20.8% 20.8%
Non-Monotone 0.0% 1.4% 0.0% 0.0% 0.4% 0.4%
Non-Convex 0.0% 1.4% 0.0% 1.5% 2.3% 1.9%
Ill-behaved 0.0% 2.8% 0.0% 1.5% 2.6% 1.9%
Peaking 0.0% 0.0% 0.0% 0.4% 0.4% 0.4%
Dipping 0.0% 0.0% 0.0% 0.0% 0.0% 0.0%
Table 8: Statistics of 24 learners in LCDB 1.1.
Learner / Ratio(%) LCDB 1.1 CC-18 (72) no FS LCDB 1.1 CC-18 (72) min-max FS LCDB 1.1 CC-18 (72) standardization FS
Missing Flat Non-Monotone Non-Convex Ill-behaved Peaking Dipping Missing Flat Non-Monotone Non-Convex Ill-behaved Peaking Dipping Missing Flat Non-Monotone Non-Convex Ill-behaved Peaking Dipping
SVM_Linear 0.0 0.0 4.2 4.2 5.6 2.8 2.8 0.0 5.6 1.4 0.0 1.4 0.0 0.0 0.0 1.4 0.0 1.4 1.4 0.0 0.0
SVM_Poly 0.0 19.4 0.0 4.2 4.2 2.8 0.0 0.0 2.8 0.0 1.4 1.4 0.0 0.0 0.0 6.9 0.0 4.2 4.2 0.0 0.0
SVM_RBF 0.0 19.4 1.4 11.1 11.1 6.9 1.4 0.0 15.3 0.0 8.3 8.3 2.8 1.4 0.0 12.5 0.0 4.2 4.2 0.0 0.0
SVM_Sigmoid 0.0 11.1 61.1 33.3 65.3 2.8 56.9 0.0 11.1 58.3 33.3 62.5 4.2 56.9 0.0 6.9 44.4 23.6 47.2 0.0 54.2
Decision Tree 0.0 2.8 0.0 0.0 0.0 0.0 1.4 0.0 0.0 1.4 0.0 1.4 0.0 1.4 0.0 0.0 1.4 0.0 1.4 0.0 1.4
ExtraTree 0.0 2.8 1.4 0.0 1.4 0.0 0.0 0.0 0.0 1.4 0.0 1.4 0.0 0.0 0.0 0.0 1.4 0.0 1.4 0.0 0.0
LogisticRegression 0.0 2.8 0.0 4.2 4.2 0.0 0.0 0.0 13.9 0.0 0.0 0.0 0.0 2.8 0.0 4.2 0.0 2.8 2.8 0.0 0.0
PassiveAggressive 0.0 1.4 5.6 0.0 5.6 0.0 1.4 0.0 1.4 0.0 0.0 0.0 0.0 0.0 0.0 0.0 1.4 0.0 1.4 0.0 0.0
Perceptron 0.0 0.0 4.2 1.4 4.2 1.4 1.4 0.0 0.0 0.0 0.0 0.0 0.0 1.4 0.0 0.0 1.4 1.4 1.4 0.0 1.4
RidgeClassifier 0.0 5.6 18.1 20.8 20.8 13.9 4.2 0.0 11.1 1.4 5.6 5.6 1.4 1.4 0.0 4.2 16.7 18.1 18.1 12.5 1.4
SGDClassifier 0.0 0.0 2.8 0.0 2.8 0.0 2.8 0.0 0.0 0.0 0.0 0.0 0.0 1.4 0.0 0.0 2.8 2.8 4.2 0.0 2.8
MLP 0.0 2.8 18.1 26.4 29.2 8.3 2.8 0.0 6.9 16.7 18.1 19.4 6.9 1.4 0.0 2.8 2.8 5.6 6.9 0.0 0.0
LDA 0.0 2.8 43.1 47.2 48.6 37.5 1.4 0.0 2.8 43.1 47.2 48.6 37.5 1.4 0.0 1.4 44.4 48.6 51.4 37.5 1.4
QDA 0.0 2.8 44.4 40.3 52.8 12.5 30.6 0.0 2.8 47.2 50.0 59.7 13.9 33.3 0.0 0.0 47.2 48.6 58.3 9.7 31.9
BernoulliNB 0.0 19.4 13.9 19.4 26.4 8.3 11.1 0.0 13.9 18.1 9.7 20.8 0.0 19.4 0.0 6.9 4.2 5.6 6.9 0.0 5.6
MultinomialNB 22.2 2.8 1.4 2.8 4.2 1.4 4.2 0.0 12.5 6.9 5.6 12.5 1.4 9.7 100.0 0.0 0.0 0.0 0.0 0.0 0.0
ComplementNB 22.2 2.8 6.9 2.8 8.3 0.0 5.6 0.0 1.4 15.3 11.1 18.1 1.4 15.3 100.0 0.0 0.0 0.0 0.0 0.0 0.0
GaussianNB 0.0 2.8 18.1 13.9 20.8 2.8 19.4 0.0 2.8 23.6 19.4 25.0 6.9 22.2 0.0 1.4 26.4 22.2 29.2 6.9 25.0
KNN 0.0 13.9 4.2 2.8 5.6 0.0 8.3 0.0 11.1 2.8 2.8 4.2 0.0 2.8 0.0 12.5 6.9 2.8 6.9 0.0 5.6
NearestCentroid 0.0 4.2 4.2 2.8 5.6 0.0 5.6 0.0 4.2 18.1 4.2 18.1 0.0 19.4 0.0 2.8 6.9 1.4 6.9 0.0 11.1
ens.ExtraTrees 0.0 8.3 1.4 0.0 1.4 0.0 1.4 0.0 6.9 1.4 0.0 1.4 0.0 1.4 0.0 6.9 1.4 0.0 1.4 0.0 1.4
ens.RandomForest 0.0 6.9 1.4 0.0 1.4 0.0 1.4 0.0 6.9 1.4 0.0 1.4 0.0 1.4 0.0 6.9 1.4 0.0 1.4 0.0 1.4
ens.GradientBoosting 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
DummyClassifier 0.0 73.6 15.3 65.3 69.4 62.5 2.8 0.0 73.6 15.3 65.3 69.4 62.5 2.8 0.0 73.6 15.3 65.3 69.4 62.5 2.8
Learner / Ratio(%) LCDB 1.1 FULL (265) no FS LCDB 1.1 FULL (265) min-max FS LCDB 1.1 FULL (265) standardization FS
Missing Flat Non-Monotone Non-Convex Ill-behaved Peaking Dipping Missing Flat Non-Monotone Non-Convex Ill-behaved Peaking Dipping Missing Flat Non-Monotone Non-Convex Ill-behaved Peaking Dipping
SVM_Linear 0.0 3.4 3.8 4.5 5.7 2.3 2.6 0.0 7.9 2.6 2.3 3.4 1.1 1.1 0.0 2.3 3.4 4.2 5.3 1.9 2.3
SVM_Poly 0.0 17.0 3.4 6.4 7.9 3.8 3.4 0.0 5.7 1.5 2.6 3.8 0.8 1.9 0.0 12.8 2.3 5.3 6.0 2.3 1.9
SVM_RBF 0.0 18.1 3.8 15.5 16.2 7.2 2.3 0.0 13.2 3.0 9.8 11.3 4.2 1.9 0.0 12.1 3.4 6.0 7.2 1.5 0.8
SVM_Sigmoid 0.0 19.2 48.3 37.7 58.5 7.9 46.4 0.0 20.0 44.5 34.7 55.8 8.7 45.7 0.0 11.7 39.6 24.9 42.6 0.8 42.6
Decision Tree 0.0 4.5 0.8 1.5 1.5 0.8 1.5 0.0 2.3 1.5 1.5 2.3 0.8 1.5 0.0 3.0 1.5 1.5 2.3 0.8 1.5
ExtraTree 0.0 3.8 1.9 0.4 1.9 0.0 1.1 0.0 2.3 1.9 0.4 1.9 0.0 1.1 0.0 2.6 1.9 0.4 1.9 0.0 1.1
LogisticRegression 0.0 6.8 2.3 4.5 5.3 1.5 1.9 0.0 11.3 1.5 1.5 1.9 1.5 2.3 0.0 7.9 3.4 4.2 4.9 3.0 1.5
PassiveAggressive 0.0 5.7 8.3 4.9 9.4 2.3 3.8 0.0 5.3 2.3 0.8 2.6 0.4 1.9 0.0 1.9 4.2 1.9 4.9 1.1 3.0
Perceptron 0.0 3.0 3.0 2.3 3.8 1.1 1.9 0.0 1.9 1.5 0.8 1.9 0.8 1.1 0.0 1.1 3.0 1.9 3.0 1.1 2.3
RidgeClassifier 0.0 7.2 14.7 15.1 17.0 10.6 5.3 0.0 10.9 2.3 3.8 4.2 1.9 1.9 0.0 6.0 14.7 15.8 17.7 10.2 2.6
SGDClassifier 0.0 2.3 3.0 1.1 3.4 0.8 3.0 0.0 3.0 1.1 1.1 1.5 1.1 2.3 0.0 0.8 3.4 3.0 4.2 1.5 1.9
MLP 0.0 4.9 21.1 21.9 27.9 8.3 4.2 0.0 7.5 18.5 23.0 26.8 6.4 2.3 0.0 3.0 9.4 14.0 16.6 3.0 1.9
LDA 0.0 3.8 32.5 33.2 37.7 24.5 7.2 0.0 4.2 32.5 33.2 37.7 24.5 7.2 0.0 3.0 32.1 33.6 37.7 23.8 7.2
QDA 0.0 3.8 34.0 37.4 46.0 12.1 27.5 0.0 3.4 35.1 40.8 48.7 11.7 30.6 0.0 1.9 38.1 40.0 47.5 9.8 34.0
BernoulliNB 0.0 26.4 12.1 23.4 28.7 13.6 8.3 0.0 22.3 17.7 12.5 23.4 2.6 19.2 0.0 9.1 8.3 8.3 12.1 1.1 9.4
MultinomialNB 30.2 9.1 4.9 7.2 8.3 2.3 6.0 0.0 14.3 14.3 9.8 18.9 2.6 19.6 100.0 0.0 0.0 0.0 0.0 0.0 0.0
ComplementNB 30.2 8.3 6.8 5.7 8.3 1.5 7.5 0.0 5.3 17.7 11.7 19.2 2.6 18.5 100.0 0.0 0.0 0.0 0.0 0.0 0.0
GaussianNB 0.0 4.5 23.0 15.8 25.3 4.5 25.7 0.0 4.5 27.5 20.0 29.4 8.7 26.4 0.0 3.4 28.3 21.1 30.6 8.3 27.5
KNN 0.0 10.9 2.6 1.9 3.8 0.8 4.5 0.0 10.2 2.6 3.0 4.5 0.8 3.0 0.0 10.9 5.7 3.4 6.8 0.8 4.5
NearestCentroid 0.0 10.9 7.5 5.3 8.3 1.1 13.2 0.0 6.8 21.5 10.2 21.9 1.5 27.2 0.0 8.7 12.8 7.5 13.2 1.1 15.5
ens.ExtraTrees 0.0 9.1 1.9 2.3 3.4 1.1 1.9 0.0 8.3 1.9 2.3 3.4 1.1 1.9 0.0 8.3 1.9 2.3 3.4 1.1 1.9
ens.RandomForest 0.0 9.1 1.5 1.9 3.0 1.1 1.5 0.0 8.7 1.5 2.3 3.0 1.1 1.1 0.0 8.3 1.5 2.3 3.0 1.1 1.1
ens.GradientBoosting 0.0 3.4 1.1 0.8 1.9 0.4 0.8 0.0 3.0 1.1 1.1 2.3 0.4 0.8 0.0 3.0 0.8 0.4 1.1 0.4 0.4
DummyClassifier 0.0 69.4 16.6 54.7 60.4 49.1 3.8 0.0 69.4 16.6 54.7 60.4 49.1 3.8 0.0 69.8 16.6 54.7 60.4 49.1 3.8
Appendix E Observation the Shape of Learning Curves
In this section, we provide additional details beyond the detection of ill-behaviors and present some results of more measurements. Specifically, we include some analyses on the localization of peakings in MLP and LDA with different feature scaling settings. Meanwhile, we visualize some results about the size of monotonicity and convexity violations (violation errors), which quantifies the severity of such behaviors. In addition, we present further learner-wise analyses related to the shape of learning curves, including the standard deviation across different random seeds. These results aim to provide deeper insights into how models behave learning differently and also show a new perspective on how our LCDB 1.1 can be used.
E.1 Surprising Shapes of MLP
In Figure 10, we show the cases where the MLP exhibits a phase transition in LCDB 1.1 CC-18. Since identifying such patterns, which characterized by abrupt improvements in performance, is somewhat subjective and they occur relatively infrequently, we did not develop a method to detect them. Instead, we selected them manually and observed that these transitions can consistently be eliminated through feature scaling.
Figure 10: The phase transition shapes of MLP in LCDB 1.1 CC-18.
In Figure 11, we visualize the location of the peak for the MLP. Specifically, the peaking detection process involves a convexity violation analysis. We identify the point of maximum convexity violation (definition in Eq. 3) and extract the coordinates of its middle point ($i^{*}$), which we take as the estimated peak position.
Figure 11: The position of peaks for the MLP for different feature scalings. Standard scaling seems to completely resolve peaking.
The location of the peaks line up suspiciously vertically. We discovered that the peak location can be explained as follows. When iterating over mini-batches, the last mini-batch can contain fewer samples because the training set sizes are not multiples of the mini-batch size. This can cause the last mini-batch to contain one or very few samples. This causes convergence issues with fitting the MLP, leading to worse performance for very specific training set sizes. Notably, the peaks also disappear when standardization scaling, which is known to improve convergence for the SGD optimizer. Thus, the peaking behavior of the MLP is largely an issue due to how mini-batches are sampled in Scikit.
E.2 Ill-Behaved Learning Curves in TabNet
The notably ill-behaved learner TabNet exhibits substantial non-convexity in its learning curves, which mostly appears to stem from phase transition phenomena. Figure 12 presents all the ill-behaved learning curves of TabNet in LCDB 1.1 CC-18. We can clearly observe that many of these curves show phase transition shapes, where the performance changes abruptly. From these observations, we find that such ill-behaviors typically occur at small training set sizes. As the training set size increases, the performance of TabNet tends to become more stable. In some cases, multiple phase transitions can even be observed within a single learning curve.
Figure 12: Ill-behaved learning curves of TabNet in LCDB 1.1 CC-18
E.3 Standard Deviation Distribution of SVM
The performance of the model improves noticeably after scaling of features (Figure 13). Furthermore, we observed substantial changes in the standard deviation over the repeats, indicating that the training process becomes more stable after feature scaling (Linear SVM as an example is shown in Figure 13).
We extract the standard deviation of all anchors from all learning curves and, based on the frequency distribution, plot three histograms: one without feature scaling, one with min-max feature scaling, and one with standardization.
Refer to caption(a)
Refer to caption(b)
Figure 13: Feature scaling improve SVM performance and can make them more stable in the context of standard deviation. (a) Comparison of SVMs best error rate in average under different feature scaling. (b) Standard deviation distribution of learner Linear SVM, the Y-axis represents the times in all anchors of the learning curve.
E.4 More Details regarding LDA
In Figure 14 we extract the location of the peak for LDA, and compare it to the dimensionality of the dataset. We see that the peak location occurs when the training set size is approximately equal to the dimensionality. Since LDA is insensitive to feature scaling, scaling the features does not lead to any changes in the learning curve (besides small differences due to numerical issues).
Refer to caption(a)
Refer to caption(b)
Figure 14: The curve shape of LDA is unaffected by feature scaling. (a) Peaking occurs when training set size is approximately equal to the dimensionality for LDA. (b) Numerical issues can cause a small change in the learning curve (this is the only curve in LCDB 1.1 that we can find that is different for LDA).
Another notable observation is that, although a substantial number of peaking behaviors in LDA learning curves are identified (see Section 5), a significant portion of peaks are still missed. Around 15% of the datasets in CC-18 contain fewer than 16 features, while the first anchor is defined at a training set size of 16. Given that $n\approx d$ in such cases, peaking behaviors are underestimated in the LDA learning curves.
To theoretically confirm that feature scaling does not affect the LDA learning curves, we provide the following proof.
Proof.
Given the LDA discriminant function:
$f_{k}(\mathbf{x})=\mathbf{x}^{T}\mathbf{w}_{k}+w_{0k},$ (6)
where
$\mathbf{w}_{k}=\bm{\Sigma}^{-1}\bm{\mu}_{k},$ (7)
$w_{0k}=-\frac{1}{2}\bm{\mu}_{k}^{T}\bm{\Sigma}^{-1}\bm{\mu}_{k}+\log p(y_{k}).$ (8)
$\mathbf{x}$ is the feature vector, $\bm{\mu}_{k}$ and $\bm{\Sigma}$ denote the class mean and shared covariance matrix, and $p(y_{k})$ is the prior probability of class $k$.
The scaling factor for feature scaling transformation $\mathbf{S}$ is a diagonal matrix, after feature scaling $\mathbf{x}^{\prime}=\mathbf{S}\mathbf{x}$. Correspondingly, the class means and covariance matrices become $\bm{\mu}_{k}^{\prime}=\mathbf{S}\bm{\mu}_{k}$ and $\bm{\Sigma}^{\prime}=\mathbf{S}\bm{\Sigma}\mathbf{S}$.
Then:
$\mathbf{w}_{k}^{\prime}$ $=(\bm{\Sigma}^{\prime})^{-1}\bm{\mu}_{k}^{\prime}$ (9)
$=(\mathbf{S}\bm{\Sigma}\mathbf{S})^{-1}(\mathbf{S}\bm{\mu}_{k})$ (10)
$=\mathbf{S}^{-1}\bm{\Sigma}^{-1}\bm{\mu}_{k}$ (11)
$=\mathbf{S}^{-1}\mathbf{w}_{k}.$ (12)
Similarly:
$w_{0k}^{\prime}$ $=-\frac{1}{2}(\bm{\mu}_{k}^{\prime})^{T}(\bm{\Sigma}^{\prime})^{-1}\bm{\mu}_{k}^{\prime}+\log p(y_{k})$ (13)
$=-\frac{1}{2}(\mathbf{S}\bm{\mu}_{k})^{T}(\mathbf{S}^{-1}\bm{\Sigma}^{-1}\mathbf{S}^{-1})\mathbf{S}\bm{\mu}_{k}+\log p(y_{k})$ (14)
$=-\frac{1}{2}\bm{\mu}_{k}^{T}\bm{\Sigma}^{-1}\bm{\mu}_{k}+\log p(y_{k})$ (15)
$=w_{0k}.$ (16)
Thus, the new discriminant function:
$f_{k}^{\prime}(\mathbf{x}^{\prime})$ $=(\mathbf{w}_{k}^{\prime})^{T}\mathbf{x}^{\prime}+w_{0k}^{\prime}$ (17)
$=(\mathbf{S}^{-1}\mathbf{w}_{k})^{T}(\mathbf{S}\mathbf{x})+w_{0k}$ (18)
$=\mathbf{w}_{k}^{T}\mathbf{x}+w_{0k}$ (19)
The new discriminant function is the same as no feature scaling one, so the learning curves should be exactly the same after feature scaling. ∎
Note that a similar argument holds for QDA, which should also be insensitive to feature scaling. However, given the reproducibility issues of QDA we did not investigate empirically.
E.5 Detailed Violations Error in CC-18
Since we define the violation error in Equations 2 and 3 to quantify the size of monotonicity and convexity violations, we are able to compare the severity of such violations across different learners and datasets. Figures 15 and 16 provide heatmap visualizations illustrating these violation errors, where zero means no violation, and white means missing learning curves.
Figure 15: Monotonicity violations error heatmap of LCDB 1.1 CC-18.
Figure 16: Convexity violations error heatmap of LCDB 1.1 CC-18.
Appendix F Learning Curves Fitting with More Parametric Models
To further validate the findings in Section 5.4, we include additional experiments with parametric models $\mathrm{MMF4}$ and $\mathrm{WBL4}$ by using the same experimental setting (Figure 17). These results confirm that the conclusions remain consistent.
Refer to caption(a)
Refer to caption(b)
Refer to caption(c)
Refer to caption(d)
Figure 17: Ill-behaved (non-monotone or non-convex) learning curves are more difficult to fit with parametric models. MSE of the fitting is proportional to the violation error. (a, b) Models: MMF4 ($\frac{ab+cx^{d}}{b+x^{d}}$). (c, d) WBL4 ($b\exp(-ax^{d})+c$).
The intrinsic properties of these parametric models conflict with the characteristics of some ill-behaved learning curves, which explains the observed experimental results. In particular, the phase transition shapes (only the convexity violated) can be effectively fitted by $\mathrm{MMF4}$. The peaking and dipping, which violate the monotonicity, cannot be modeled by these parametric models. For reference, we also provide illustrative examples (Figure 18) highlighting the alignment (or misalignment) between model properties and the geometric characteristics of learning curves, as discussed in the main text.
Refer to caption(a)
Refer to caption(b)
Refer to caption(c)
Figure 18: The parametric models fitting examples on ill-behaved shapes of learning curves. (a) Peaking. (b) Phase Transition. (c) Dipping.
Appendix G Learning Curve Crossings Affect Successive Halving
In this section, we provide detailed information about the experimental setup and results analyzing the relationship between learning curve crossings and the model selection performance of Successive Halving (SH) [61].
Figure 19 provides a pairwise crossing probability matrix for all learners to show that learning curves cross. The experiments are conducted in LCDB 1.1 CC-18 min-max feature scaling version since there are no missing curves. The left matrix shows the probability that learner A initially outperforms learner B at the lowest fidelity. The middle matrix refines this by showing the probability that A starts higher but ends lower than B, capturing the crossing from above. The right matrix shows the conditional probability of being overtaken given an early advantage, highlighting how frequently an initial lead fails to persist.
Figure 19: A simple method to evaluate crossing probability
To showcase how we can use LCDB 1.1 to study the connection between SH model selection performance and crossing of learning curves, we show more detailed experimental results for comparing the performance of SH on two groups of learners, one with rarely and one with commonly crossing learning curves. In Figure  20, the blue and orange groups are the learners whose curves cross rarely and frequently, respectively. The left, middle, and right column figures are for the case where SH starts at the first available anchor (16 training instances), the 8th anchor (30 training instances), and the 16th anchor (59 training instances), respectively. From top to bottom, each row corresponds to a different per-round budget increase rate in the SH procedure, specifically 12.5%, 25.0%, 50.0%, and 100%. The per-round budget increase rate determines how much the training budget is increased between consecutive SH rounds. The left panel of box-plots show, for different values of $k$, the probability (across the 5 outer folds of datasets) that the finally chosen algorithm is under the top $k$. However, since the final performance differences of learners may be small, we complement this figure with the regrets on the right panel of box-plots, i.e., final error rate of the chosen learner minus the minimum of final error rate (in the log-scale).
Refer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 20: Results of SH for varying starting anchors (with 16, 30, 59 training instances, respectively) with varying budgets (12.5%, 25.0%, 50.0%, and 100%).
The figures nicely confirm the intuition on the relevance of learning curve crossing and the performance of SH. In the group of learners whose curves rarely cross (blue), the algorithm almost always picks the best or at least runner-up. For learners that frequently cross this is less often the case. The regrets show a similar pattern. As such, we can observe that crossing curves make model selection using SH significantly more challenging. The difference between the groups is also nicely reflected in the regrets, which are significantly better in the group of learners whose curves rarely cross compared to the ones where curves frequently cross.
When SH starts at the first available anchor (16 training instances), its ability to identify top-performing algorithms is particularly poor for the frequently-crossing group: in approximately 80% of cases, it fails to select either the best or the second-best candidate. This suggests that critical curve crossings likely occur at very early stages, causing premature elimination of ultimately superior learners.
However, this issue diminishes as the starting anchor increases. When starting at the 16th anchor (59 training instances), the performance gap narrows, and SH becomes more effective even for the frequently-crossing group. These results demonstrate that crossing learning curves pose a serious challenge for multi-fidelity optimization strategies like SH, especially when early budgets dominate the selection process.
Appendix H Alternative Method To Detect Monotonicity Violations
Monotonicity can also be assessed at the local level by examining trends between consecutive anchors. We introduce a method to statistically identify local monotonicity, classifying all segments between consecutive anchor pairs into three categories: significant improvement, significant worsening, and insignificant change. This method evaluates all consecutive segments of the learning curve and classifies each segment into one of three categories: improvement, worsening, or insignificant, based on statistical significance of paired t-test (an example in Figure 21).
By leveraging the local monotonicity of consecutive anchors, we propose an alternative approach to detect the occurrence of the peaking phenomenon. Specifically, we assume that there is always at least a peaking occurrence between a pair of improvement and worsening segments, potentially interspersed with insignificant status in between. Based on this, we define a criterion for detecting peaks by examining such anchor status transitions across the learning curve. The resulting detection, illustrated in Figure 21, provides a complementary perspective to the main method. Although this approach does not employ the Bonferroni correction and is therefore slightly more permissive, the estimated probabilities of peak occurrences remain broadly consistent with Figure 4.
However, due to our conservative stance toward identifying ill-behaved learning curve shapes, we opt for a more statistically rigorous approach. We do not adopt the peak detection method by using local monotonicity described above, as it does not incorporate multiple comparison corrections and may be prone to false positives.
Refer to caption(a)
Refer to caption(b)
Figure 21: Illustration of (a) local monotonicity identification and (b) the results of using local monotonicity to detect statistic peaking.
Appendix I Statistical Correction Using Holm’s Method
We consider to use a slightly less conservative method called Holm’s Step-Down Procedure (Holm’s method) [18], in the sense that it will reject more null hypotheses, typically resulting in fewer Type II errors. As shown in Table 9, the results remain consistent with our original findings: an even larger fraction of learning curves are identified as ill-behaved, up to 19%.
Table 9: Ill-behavior statistics of the LCDB 1.1 (23 learners, Dummy excluded) by using Holm’s method.
Shapes / Database LCDB 1.1 CC-18 (72) LCDB 1.1 FULL (265)
no FS min-max FS standardization FS no FS min-max FS standardization FS
Missing 1.9% 0.0% 8.7% 2.6% 0.0% 8.7%
Non-Monotone ($\neg$ M) 13.3% 13.0% 10.6% 13.2% 14.1% 11.8%
Non-Convex ($\neg$ C) 14.6% 14.1% 11.8% 14.8% 15.4% 13.5%
Ill-behaved ($\neg$ M $\cup$ $\neg$ C) 18.2% 17.3% 14.7% 18.8% 19.8% 16.8%
Appendix J Broader Discussion about LCDB 1.1
Resource Usage and Green Machine Learning.
The creation of LCDB 1.1 involved approximately 800,000 CPU-hours and 3,000 GPU-hours of computing. The computing was conducted on a heterogeneous cluster environment provided by the Delft Artificial Intelligence Cluster (DAIC) [77]. Since the specific CPU node used for each job was not fixed, jobs were scheduled across a range of CPU nodes: AMD EPYC 7502P 32-Core Processor, AMD EPYC 9534 64-Core Processor, AMD EPYC 7413 24-Core Processor, AMD EPYC 7452 32-Core Processor, AMD EPYC 7543 32-Core Processor, Intel(R) Xeon(R) CPU E5-2667 v4 @ 3.20GHz, Intel(R) Xeon(R) CPU E5-2680 v4 @ 2.40GHz, Intel(R) Xeon(R) Gold 5218 CPU @ 2.30GHz, Intel(R) Xeon(R) Gold 6130 CPU @ 2.10GHz, Intel(R) Xeon(R) Gold 6140 CPU @ 2.30GHz. For TabPFN v2 learning curves, we used DelftBlue Supercomputer hosted at the Delft High Performance Computing Centre (DHPC) [78], utilizing Nvidia A100 and Nvidia Tesla V100. Each compute job was allocated 20 GB of memory. We set the maximum execution time as 3.5 hours for computing per anchor, inner, and outer seed combination. If the total runtime exceeds this limit, the job is terminated and its status is recorded as a timeout. Both the error message and execution status were logged.
We acknowledge the environmental footprint of this computation and have taken steps to reuse existing model outputs where possible to minimize redundant training. All prediction results during our training process are stored, which allows us to obtain more variants in addition to error rate learning curves, such as the AUC curve, the F1 score learning curve, the log-loss learning curve, and some more types of learning curves in metrics that may be of interest in the future. Furthermore, the processed learning curve data, LCDB 1.1, are annotated with standardized metadata using the Croissant format [79], which facilitates dataset discoverability and machine-readability through integration with web-based data indexing systems.
Social Impact.
LCDB 1.1 provides researchers with a perspective from learning curve to better understand the relationship between model performance and the amount of training data. This can be especially valuable in domains where data collection is costly or limited, such as medicine [80], potentially improving outcomes in areas with high social impact. Furthermore, from a meta-learning standpoint, LCDB 1.1 could be a challenging benchmark contributing toward more efficient and automated machine learning systems, which can democratize access to high-quality models in fields that traditionally require significant expert knowledge. However, as with many powerful tools, LCDB 1.1 also poses dual-use concerns. Insights derived from learning curves and dataset performance may inadvertently aid malicious applications such as targeted misinformation or the development of weapons systems.
Appendix K OpenML Dataset List in LCDB 1.1
Since data curation is far from trivial, even for tabular data [81], we show the following basic properties of the dataset used in LCDB 1.1: OpenML ID (ID), name of the dataset (Name), number of features (#Features), number of samples (#Samples), number of classes (#Classes), and the maximum class ratio (Ratio). Hopefully, this summary can facilitate an assessment of the composition of LCDB 1.1 and allows for the identification of relevant subsets for further analysis.
Table 10: A total overview of 265 OpenML datasets used in LCDB 1.1 FULL.
ID Name #Features #Samples #Classes Ratio
3 kr-vs-kp 36 3196 2 0.52
6 letter 16 20000 26 0.04
11 balance-scale 4 625 3 0.46
12 mfeat-factors 216 2000 10 0.10
13 breast-cancer 9 286 2 0.70
14 mfeat-fourier 76 2000 10 0.10
15 breast-w 9 699 2 0.66
16 mfeat-karhunen 64 2000 10 0.10
18 mfeat-morphological 6 2000 10 0.10
21 car 6 1728 4 0.70
22 mfeat-zernike 47 2000 10 0.10
23 cmc 9 1473 3 0.43
24 mushroom 22 8124 2 0.52
26 nursery 8 12960 5 0.33
28 optdigits 64 5620 10 0.10
29 credit-approval 15 690 2 0.56
30 page-blocks 10 5473 5 0.90
31 credit-g 20 1000 2 0.70
32 pendigits 16 10992 10 0.10
36 segment 19 2310 7 0.14
37 diabetes 8 768 2 0.65
38 sick 29 3772 2 0.94
44 spambase 57 4601 2 0.61
46 splice 60 3190 3 0.52
50 tic-tac-toe 9 958 2 0.65
54 vehicle 18 846 4 0.26
55 hepatitis 19 155 2 0.79
57 hypothyroid 29 3772 4 0.92
60 waveform-5000 40 5000 3 0.34
61 iris 4 150 3 0.33
151 electricity 8 45312 2 0.58
179 adult 14 48842 2 0.76
180 covertype 54 110393 7 0.47
181 yeast 8 1484 10 0.31
182 satimage 36 6430 6 0.24
184 kropt 6 28056 18 0.16
185 baseball 16 1340 3 0.91
188 eucalyptus 19 736 5 0.29
201 pol 48 15000 11 0.62
273 IMDB.drama 1001 120919 2 0.64
293 covertype 54 581012 2 0.51
299 libras_move 90 360 15 0.07
300 isolet 617 7797 26 0.04
307 vowel 12 990 11 0.09
336 SPECT 22 267 2 0.79
346 aids 4 50 2 0.50
351 codrna 8 488565 2 0.67
354 poker 10 1025010 2 0.50
357 vehicle_sensIT 100 98528 2 0.50
380 SyskillWebert-Bands 2 61 3 0.64
389 fbis.wc 2000 2463 17 0.21
390 new3s.wc 26832 9558 44 0.07
391 re0.wc 2886 1504 13 0.40
392 oh0.wc 3182 1003 10 0.19
393 la2s.wc 12432 3075 6 0.29
395 re1.wc 3758 1657 25 0.22
396 la1s.wc 13195 3204 6 0.29
398 wap.wc 8460 1560 20 0.22
399 ohscal.wc 11465 11162 10 0.15
401 oh10.wc 3238 1050 10 0.16
446 prnn_crabs 7 200 2 0.50
458 analcatdata_authorship 70 841 4 0.38
469 analcatdata_dmft 4 797 6 0.19
554 mnist_784 784 70000 10 0.11
679 rmftsa_sleepdata 2 1024 4 0.39
715 fri_c3_1000_25 25 1000 2 0.56
718 fri_c4_1000_100 100 1000 2 0.56
720 abalone 8 4177 2 0.50
722 pol 48 15000 2 0.66
723 fri_c4_1000_25 25 1000 2 0.55
727 2dplanes 10 40768 2 0.50
728 analcatdata_supreme 7 4052 2 0.76
734 ailerons 40 13750 2 0.58
735 cpu_small 12 8192 2 0.70
737 space_ga 6 3107 2 0.50
740 fri_c3_1000_10 10 1000 2 0.56
741 rmftsa_sleepdata 2 1024 2 0.50
743 fri_c1_1000_5 5 1000 2 0.54
751 fri_c4_1000_10 10 1000 2 0.56
752 puma32H 32 8192 2 0.50
761 cpu_act 21 8192 2 0.70
772 quake 3 2178 2 0.56
797 fri_c4_1000_50 50 1000 2 0.56
799 fri_c0_1000_5 5 1000 2 0.50
803 delta_ailerons 5 7129 2 0.53
806 fri_c3_1000_50 50 1000 2 0.56
807 kin8nm 8 8192 2 0.51
813 fri_c3_1000_5 5 1000 2 0.56
816 puma8NH 8 8192 2 0.50
819 delta_elevators 6 9517 2 0.50
821 house_16H 16 22784 2 0.70
822 cal_housing 8 20640 2 0.59
823 houses 8 20640 2 0.57
833 bank32nh 32 8192 2 0.69
837 fri_c1_1000_50 50 1000 2 0.55
843 house_8L 8 22784 2 0.70
845 fri_c0_1000_10 10 1000 2 0.51
846 elevators 18 16599 2 0.69
847 wind 14 6574 2 0.53
849 fri_c0_1000_25 25 1000 2 0.50
866 fri_c2_1000_50 50 1000 2 0.58
871 pollen 5 3848 2 0.50
881 mv 10 40768 2 0.60
897 colleges_aaup 15 1161 2 0.70
901 fried 10 40768 2 0.50
903 fri_c2_1000_25 25 1000 2 0.56
904 fri_c0_1000_50 50 1000 2 0.51
910 fri_c1_1000_10 10 1000 2 0.56
912 fri_c2_1000_5 5 1000 2 0.58
913 fri_c2_1000_10 10 1000 2 0.58
914 balloon 1 2001 2 0.76
917 fri_c1_1000_25 25 1000 2 0.55
923 visualizing_soil 4 8641 2 0.55
930 colleges_usnews 33 1302 2 0.53
934 socmob 5 1156 2 0.78
953 splice 60 3190 2 0.52
958 segment 19 2310 2 0.86
959 nursery 8 12960 2 0.67
962 mfeat-morphological 6 2000 2 0.90
966 analcatdata_halloffame 16 1340 2 0.91
971 mfeat-fourier 76 2000 2 0.90
976 JapaneseVowels 14 9961 2 0.84
977 letter 16 20000 2 0.96
978 mfeat-factors 216 2000 2 0.90
979 waveform-5000 40 5000 2 0.66
980 optdigits 64 5620 2 0.90
991 car 6 1728 2 0.70
993 kdd_ipums_la_97-small 60 7019 2 0.63
995 mfeat-zernike 47 2000 2 0.90
1000 hypothyroid 29 3772 2 0.92
1002 ipums_la_98-small 55 7485 2 0.89
1018 ipums_la_99-small 56 8844 2 0.94
1019 pendigits 16 10992 2 0.90
1020 mfeat-karhunen 64 2000 2 0.90
1021 page-blocks 10 5473 2 0.90
1036 sylva_agnostic 216 14395 2 0.94
1040 sylva_prior 108 14395 2 0.94
1041 gina_prior2 784 3468 10 0.11
1042 gina_prior 784 3468 2 0.51
1049 pc4 37 1458 2 0.88
1050 pc3 37 1563 2 0.90
1053 jm1 21 10885 2 0.81
1056 mc1 38 9466 2 0.99
1063 kc2 21 522 2 0.80
1067 kc1 21 2109 2 0.85
1068 pc1 21 1109 2 0.93
1069 pc2 36 5589 2 1.00
1083 mouseType 45101 214 7 0.32
1084 BurkittLymphoma 22283 220 3 0.58
1085 anthracyclineTaxaneChemotherapy 61359 159 2 0.60
1086 ovarianTumour 54621 283 3 0.87
1087 hepatitisC 54621 283 3 0.87
1088 variousCancers_final 54675 383 10 0.40
1116 musk 167 6598 2 0.85
1119 adult-census 14 32561 2 0.76
1120 MagicTelescope 10 19020 2 0.65
1128 OVA_Breast 10935 1545 2 0.78
1130 OVA_Lung 10935 1545 2 0.92
1134 OVA_Kidney 10935 1545 2 0.83
1138 OVA_Uterus 10935 1545 2 0.92
1139 OVA_Omentum 10935 1545 2 0.95
1142 OVA_Endometrium 10935 1545 2 0.96
1146 OVA_Prostate 10935 1545 2 0.96
1161 OVA_Colon 10935 1545 2 0.81
1166 OVA_Ovary 10935 1545 2 0.87
1216 Click_prediction_small 9 1496391 2 0.96
1233 eating 6373 945 7 0.15
1235 Agrawal1 9 1000000 2 0.67
1236 Stagger1 3 1000000 2 0.89
1441 KungChi3 39 123 2 0.87
1448 KnuggetChase3 39 194 2 0.81
1450 MindCave2 39 125 2 0.65
1457 amazon-commerce-reviews 10000 1500 50 0.02
1461 bank-marketing 16 45211 2 0.88
1462 banknote-authentication 4 1372 2 0.56
1464 blood-transfusion-service-center 4 748 2 0.76
1465 breast-tissue 9 106 6 0.21
1468 cnae-9 856 1080 9 0.11
1475 first-order-theorem-proving 51 6118 6 0.42
1477 gas-drift-different-concentrations 129 13910 6 0.22
1478 har 561 10299 6 0.19
1479 hill-valley 100 1212 2 0.50
1480 ilpd 10 583 2 0.71
1483 ldpa 7 164860 11 0.33
1485 madelon 500 2600 2 0.50
1486 nomao 118 34465 2 0.71
1487 ozone-level-8hr 72 2534 2 0.94
1488 parkinsons 22 195 2 0.75
1489 phoneme 5 5404 2 0.71
1494 qsar-biodeg 41 1055 2 0.66
1497 wall-robot-navigation 24 5456 4 0.40
1499 seeds 7 210 3 0.33
1501 semeion 256 1593 10 0.10
1503 spoken-arabic-digit 14 263256 10 0.10
1509 walking-activity 4 149332 22 0.15
1510 wdbc 30 569 2 0.63
1515 micro-mass 1300 571 20 0.11
1566 hill-valley 100 1212 2 0.50
1567 poker-hand 10 1025009 10 0.50
1575 ijcnn 22 191681 2 0.90
1590 adult 14 48842 2 0.76
1592 aloi 128 108000 1000 0.00
1597 creditcard 29 284807 2 1.00
4134 Bioresponse 1776 3751 2 0.54
4135 Amazon_employee_access 9 32769 2 0.94
4137 Dorothea 100000 1150 2 0.90
4534 PhishingWebsites 30 11055 2 0.56
4538 GesturePhaseSegmentationProcessed 32 9873 5 0.30
4541 Diabetes130US 49 101766 3 0.54
6332 cylinder-bands 37 540 2 0.58
23381 dresses-sales 12 500 2 0.58
23512 higgs 28 98050 2 0.53
23517 numerai28.6 21 96320 2 0.51
40498 wine-quality-white 11 4898 7 0.45
40499 texture 40 5500 11 0.09
40664 car-evaluation 21 1728 4 0.70
40668 connect-4 42 67557 3 0.66
40670 dna 180 3186 3 0.52
40672 fars 29 100968 8 0.42
40677 led24 24 3200 10 0.11
40685 shuttle 9 58000 7 0.79
40687 solar-flare 12 1066 6 0.31
40701 churn 20 5000 2 0.86
40713 dis 29 3772 2 0.98
40900 Satellite 36 5100 2 0.99
40910 Speech 400 3686 2 0.98
40923 Devnagari-Script 1024 92000 46 0.02
40927 CIFAR_10 3072 60000 10 0.10
40966 MiceProtein 77 1080 8 0.14
40971 collins 19 1000 30 0.08
40975 car 6 1728 4 0.70
40978 Internet-Advertisements 1558 3279 2 0.86
40979 mfeat-pixel 240 2000 10 0.10
40981 Australian 14 690 2 0.56
40982 steel-plates-fault 27 1941 7 0.35
40983 wilt 5 4839 2 0.95
40984 segment 16 2310 7 0.14
40994 climate-model-simulation-crashes 18 540 2 0.91
40996 Fashion-MNIST 784 70000 10 0.10
41027 jungle_chess_2pcs_raw_endgame_complete 6 44819 3 0.51
41142 christine 1636 5418 2 0.50
41143 jasmine 144 2984 2 0.50
41144 madeline 259 3140 2 0.50
41145 philippine 308 5832 2 0.50
41146 sylvine 20 5124 2 0.50
41150 MiniBooNE 50 130064 2 0.72
41156 ada 48 4147 2 0.75
41157 arcene 10000 100 2 0.56
41158 gina 970 3153 2 0.51
41159 guillermo 4296 20000 2 0.60
41161 riccardo 4296 20000 2 0.75
41163 dilbert 2000 10000 5 0.20
41164 fabert 800 8237 7 0.23
41165 robert 7200 10000 10 0.10
41166 volkert 180 58310 10 0.22
41167 dionis 60 416188 355 0.01
41168 jannis 54 83733 4 0.46
41169 helena 27 65196 100 0.06
41228 Klaverjas2018 32 981541 2 0.54
41972 Indian_pines 220 9144 8 0.44
42734 okcupid-stem 19 50789 3 0.72
42742 porto-seguro 57 595212 2 0.96
42769 Higgs 28 1000000 2 0.53
42809 kits 27648 1000 2 0.52
42810 PCam 27648 4000 2 0.51
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The main claims made in the abstract and introduction accurately reflect the contributions and scope of the paper.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The paper includes a discussion section where we discuss minor flaws in the dataset and the limitations of the analysis methods we used. We also mention the possibility of further improving the dataset and the challenges of the large amount of computing resources required.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: This is a dataset paper, so most of the content is centered around empirical findings. The only theoretical hypothesis is proved in the appendix.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We have tried our best to make the LCDB 1.1 fully reproducible, by using a docker container and fixed python package versions, yet a few QDA curves and 1 LDA curve are non-reproducible due to the Scikit implementation. We have provided details regarding this.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We provide open access to the data and all the code with detailed instruction for reproducibility.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We clearly state the adequate experimental setup in the main text and some parts provide more details in the appendices. The provided code includes all the random seeds (initialization and splits), and we include a docker image and package versions.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: Depending on the situation, we report either error bars or statistical significance to ensure the robustness of our findings.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We put the computing hours and discussion about green machine learning in Appendix J. For the analysis part, the computational time is negligible (less than 3 hours of CPU time).
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research adheres to the NeurIPS Code of Ethics. The dataset was collected and processed in accordance with ethical standards.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We have a short but meaningful discussion about social impact in Appendix J.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Our dataset is about the performance of different learning algorithms and the relationship between the required training data. We think there is no safeguard risk.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The dataset used in this work is constructed from publicly available data sources hosted on OpenML. These datasets are released under the CC BY 4.0 license, which permits use, modification, and redistribution with appropriate attribution.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We release the database under the CC BY 4.0 license. We provide all the scripts to reproduce the figures of the paper, illustrating the usage of the LCDB 1.1. We also provide the demo notebook and readme in the GitHub repository.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The research does not involve human subjects or crowdsourcing, and thus no Institutional Review Board (IRB) approval or equivalent was required.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: LLM is used only for writing, editing, or formatting purposes.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.